OpenStudy (anonymous):

Need help with these questions please! Simplify (x - y + 1) - (x + y - 1). 2x - 2y + 2 -2y + 2 2y - 2 -2x + 2y Given: Q = 7m + 3n, R = 11 - 2m, S = n + 5, and T = -m - 3n + 8. Simplify [Q - R] + [S - T]. 10m + 5n - 24 10m - 5n + 24 10m + 7n - 14 10m - 7n - 14 Simplify {n - 1 - [n - 1 - (n - 1)]}. n - 1 -n - 1 3n - 3 -3n + 3 Given: Q = 7m + 3n, R = 11 - 2m, S = n + 5, and T = -m - 3n + 8. Simplify R - S + T. m - 2n - 2 -m + 2n - 2 3m - 4n + 14 -3m - 4n + 14 I only need the answer but if you could tell me how you got the answer I would appreciate it!

OpenStudy (anonymous):

Simple math first you got to do Simplify {n - 1 - [n - 1 - (n - 1)]}. then you have to Simplify R - S + T. m - 2n - 2 -m + 2n - 2 3m - 4n + 14 -3m - 4n + 14 Last but not least you do this Given: Q = 7m + 3n, R = 11 - 2m, S = n + 5, and T = -m - 3n + 8. Simplify [Q - R] + [S - T]. 10m + 5n - 24 10m - 5n + 24 10m + 7n - 14 10m - 7n - 14

OpenStudy (anonymous):

1.y = -x -1 y = -(-1) -1 = +1-1 = 0 2.3n+5=7(n-7) 3.(1/2!)+(2/3!)+(3/4 4.q+(q^2-4pr)^1/2